AHOPKEEPER HAS THREE TYPES OF BLACK PULSE COSTING RS.60, RS.72 AND RS.90 PER KG RESPECTIVELY. OUT OF THEM THE FIRST TWO TYPES OF PULSE ARE MIXED IN THE RATION OF 2:3. IN WHAT RATIO THE THINRD TYPE OF PULSE SHOULD BE MIXED SO THAT THE COST OF MIXTURE WILL BE RS.78 PER KG?
Answer by palanisamy(496)

A SHOPKEEPER HAS THREE TYPES OF BLACK PULSE COSTING RS.60, RS.72 AND RS.90 PER KG RESPECTIVELY.
THE FIRST TWO TYPES OF PULSE ARE MIXED IN THE RATION OF 2:3.
lET THE THIRD TYPE PULSE BE MIXED IN THE RATIO x
THE FINAL COST OF THE MIXTURE IS RS.78 PER KG
THEREFORE, WE GET
(2X60)+(3X72)+90x = (2+3+x)X78
120+216+90x = (5+x)X 78
336+90x = 390+78x
90x-78x = 390-336
12x = 54
x = 54/12
x = 9/2